# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ Acknowledgements
The author would like to thank many people for useful discussions on this project, including Kilian Bönisch, Steven Charlton, Chi Hong Chow, Henri Cohen, Håvard Damm-Johnsen, Vasily Golyshev, Wen-Ching Winnie Li, Yingkun Li, Ling Long, Danylo Radchenko, Berend Ringeling, Fernando Rodriguez Villegas, Hasan Saad, and Jan Vonk. The author is in particular grateful to Don Zagier and Wadim Zudilin for all the useful ideas on proofs and generalizations.
This project originated from discussions during the International Conference on Modular Forms and $q$-Series at University of Cologne, and during the follow-up workshop at the Hausdorff Research Institute for Mathematics to the trimester program “Periods in Number Theory, Algebraic Geometry and Physics” funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. Part of the later discussions also took place during the author’s visit to the Abdus Salam International Centre for Theoretical Physics hosted by Vasily Golyshev and Don Zagier, and the author’s visit to Louisiana State University hosted by Ling Long and partially supported by the Simons Foundation grant MP-TSM-00002492.
@@ 2. Simple poles
Conjecture 2.1.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
(2) $a_{np}(F_{k,C})\;\equiv\;a_{p}(C)^{k-2}a_{n}(F_{k,C})\pmod{p}$.
Remark 2.2.
2.1 is a theorem if we further assume that $p\geq 5$, from which Theorem 1.3 follows. In fact, the cases when $k\in\{4,6,8,10,14\}$ all follow from the case when $k=4$, and the proof when $k=4$ depends on the aforementioned connection with hypergeometric functions, namely the congruence
$a_{p}(F_{4,C})$ $\;\equiv\;\big(j(C)(j(C)-1728)\big)^{\frac{p-1}{2}}\cdot\sum_{m=0}^{p-1}\frac{(6m)!}{(m!)^{3}(3m)!}\cdot j(C)^{-m}\pmod{p}$.
We will show all above in Section 5, where we will prove a suitable version of 2.1 for elliptic curves over general number fields.
From a different point of view, 2.1 describes the behavior of $F_{k,C}$ under the $U_{p}$ operator. In particular, it says that $F_{k,C}$, when viewed as a mod $p$ modular form, is a $U_{p}$-eigenform with eigenvalue $a_{p}(C)^{k-2}$. In general, one can also describe the behavior of $F_{k,C}$ under the $U_{p}^{l}$ operator if one considers the supersingular and the ordinary cases separately.
@@
While Remark 2.2 already presents a connection between meromorphic modular forms and hypergeometric functions, the type of congruences that appear in 2.3 and 2.4 may also remind one of hypergeometric (super)congruences.
In the case of CM elliptic curves, it is possible to unify the supersingular and ordinary dichotomy via a single, but slightly weaker, congruence, in the form of the three-term ASD congruence as mentioned in Section 1. There are two viewpoints in the CM setting: one can interpret everything in terms of CM elliptic curves or in terms of negative discriminants. We will elaborate on both viewpoints here.
@@
(6) $a_{np^{l}}(F_{k,C})-a_{p}(\Theta){\hskip 1.0pt}a_{np^{l-1}}(F_{k,C})+\chi_{K}(p)p^{k-2}a_{np^{l-2}}(F_{k,C})\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
@@
By Deuring’s criterion (see [Deu41] or [Sil94, Chapter II, Exercise 2.30]), a good prime $p$ of $C$ is ordinary (resp. supersingular) if and only if $\chi_{K}(p)=1$, i.e. $p$ splits in $K$ (resp. $\chi_{K}(p)=-1$, i.e., $p$ is inert in $K$). The coefficient $a_{p}(\Theta)$ for a good prime $p$ of $C$ can then be written down explicitly as
@@
where $\alpha_{p}$ and $\beta_{p}$ are the two roots of $X^{2}-a_{p}(C)X+p$.
From this explicit description, one can readily deduce 2.6 from the supersingular (2.3) and ordinary (2.4) dichotomy. In the supersingular case, i.e., $\chi_{K}(p)=-1$, 2.6 reduces to that
@@
which follows from 2.3. In the ordinary case, i.e., $\chi_{K}(p)=1$, 2.6 reduces to that
@@
$\bigg(a_{np^{l}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\bigg)$
$\hskip 42.67912pt+\bigg(\frac{p}{u_{p}(C)}\bigg)^{k-2}\bigg(a_{np^{l-1}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-2}}(F_{k,C})\bigg)\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
which follows from 2.4.
Now, we turn to the negative discriminants. Let $D<0$ (with $D\neq-3,-4$) be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $\alpha_{D}\in\mathcal{H}$ be a fixed CM point of discriminant $D$. For $k\in\{4,6,8,10,14\}$, write
@@
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form (in fact Hecke eigenform)
@@
One can also write 2.4 in the CM case explicitly in terms of negative discriminants.
@@
Let $k\in\{4,6,8,10,14\}$ and let $D<0$ be a discriminant with class number $1$. Let $p$ be a prime with $\big(\frac{D}{p}\big)=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$, and write $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,D})$ $\;\equiv\;\pi^{k-2}a_{np^{l-1}}(F_{k,D})\pmod{\overline{\pi}^{(k-1)l}}$
$a_{np^{l}}(F_{k,D})$ $\;\equiv\;\overline{\pi}^{k-2}a_{np^{l-1}}(F_{k,D})\pmod{\pi^{(k-1)l}}$.
@@ 3. Poles of higher order
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}$
@@
The consideration of this particular space is motivated by several reasons. It is easy to check that for $r\geq k$, $\frac{E_{k}}{(j-j(C))^{r}}$ can be written as a linear combination of $\frac{E_{k}}{j-j(C)},\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}$, and $\mathcal{D}^{k-1}f$ for some modular form $f$ of weight 2k2-k, where $\mathcal{D}=q\frac{d}{dq}$ is the usual derivative. From the viewpoint of the magnetic property, the modular form $\mathcal{D}^{k-1}f$ here is trivially (k1)(k-1)-magnetic, so it suffices to consider the denominator only up to the (k1)(k-1)-st power. More importantly, the discussion above yields a non-canonical isomorphism (for $k\in\{4,6,8,10,14\}$ and $j(C)\notin\{0,1728\}$)
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}\;\cong\;\frac{S_{k}^{\mathrm{mero},C}}{\mathcal{D}^{k-1}M_{2-k}^{\mathrm{mero},C}},$
where $M_{k}^{\mathrm{mero},C}$ denotes the space of meromorphic modular forms of level $1$ and weight $k$ with precisely one pole at (the point on $\mathcal{H}$ corresponding to) $C$, and $S_{k}^{\mathrm{mero},C}$ denotes its subspace consisting of elements having vanishing constant terms in the Fourier expansions. In Kazalicki–Scholl [KS16], the similar quotient where $C$ is replaced by the cusp $\infty$ already occurred and was used to prove the ASD congruences satisfied by weakly holomorphic modular forms. Specifically, they provided a cohomological interpretation of the quotient and transferred the ASD congruences to the action of the Frobenius at $p$ on the cohomology side. In our case, a cohomological interpretation of the quotient also exists and was discussed in Brown–Fonseca [BF25], where they also related it to $\mathrm{Sym}^{k-2}C$. This suggests that these meromorphic modular forms should be linked to symmetric powers of elliptic curves through the ASD congruences.
The guiding philosophy in our case is that the action of $U_{p}$ on the modular form side  [^2] Indeed, there is no $U_{p}$ action on the space of modular forms that is considered, since the $U_{p}$ operator increases the level and shifts the pole. Still, we will formally define the $U_{p}$ operator by $\sum a_{n}q^{n}\mapsto\sum a_{pn}q^{n}$. should correspond to the action of the Frobenius at $p$ on the elliptic curve side.
@@
where $P_{p}^{*}$ is the reciprocal polynomial of $P_{p}$, i.e., $P_{p}^{*}(X)=X^{k-1}P_{p}(X^{-1})$. In view of Kazalicki–Scholl [KS16], it appears that the formulation using $V_{p}$ may be the “correct” one, since their proof of the ASD congruences involves identifying a $V_{p}$-related operator with the Frobenius at $p$. It may thus be the case that we should also compare the $V_{p}$ operator on the modular form side with the Frobenius at $p$ on the elliptic curve side in our case.
If one separates the supersingular and ordinary primes as before, then one also observes a generalization of 2.3. However, generalizations of 2.4 seem to only exist in the CM case, which will be discussed in Section 4.
@@
When $p$ is supersingular, $F_{k,C}^{(r)}$ may be viewed as an overconvergent $p$-adic modular form, since the pole lies in the supersingular locus. Indeed, for $l\geq 2$, 3.3 leads to the consideration of $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$. Here one observes that there is no old cusp form by our choice of $k$ and the operator $U_{p}-p^{k-2}$ annihilates all the $p$-new cusp forms of weight $k$. Hence, heuristically one should be able to write $(U_{p}-p^{k-2})(F_{k,C}^{(r)})$ as an asymptotic sum of $U_{p}$-eigenforms with slopes $\geq k-1$, which then also supports the conjectural congruence satisfied by $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$. (We refer to [Cal] and [GM95] for discussions on the asymptotic expansions of overconvergent modular forms using $U_{p}$-eigenforms.)
@@
3.3 implies 3.1 in the supersingular case. We sketch a proof here for $l\geq k-1$. For supersingular $p$, the characteristic polynomial $P_{p}(X)$ is given by
@@
It then follows from 3.3 and direct computations that for all $s\in\mathbb{Z}_{\geq 0}$,
@@
We will finish this section with two comments. The first one is that one should be able to generalize the whole setting to elliptic curves over number fields. That is, one considers an elliptic curve $C$ over a number field $L$. In this case, one should then consider congruences modulo powers of $\mathfrak{p}$, and the action of the Frobenius at $\mathfrak{p}$ on the elliptic curve side should correspond to the action of $U_{\mathrm{N}(\mathfrak{p})}$ on the modular form side, where $\mathfrak{p}$ is some prime of $L$ and $\mathrm{N}(\mathfrak{p})$ denotes its norm. The second one is that the connection between meromorphic modular forms and symmetric powers of elliptic curves should be rather motivic  [^3] This is also why we always write $\mathrm{Sym}^{k-2}C$ instead of $\mathrm{Sym}^{k-2}H^{1}(C)$ with some specified cohomology, since we think $\mathrm{Sym}^{k-2}C$ as being the motive attached to the symmetric power., as discussed in Brown–Fonseca [BF25]. In particular, the congruences that we observe here should be viewed as visualizations of the motivic connection at finite places. While this is not necessarily clear to the author, it may be reasonable to expect visualizations at infinite places linking special values of $L$-functions of symmetric powers of elliptic curves to “periods” of meromorphic modular forms (see also [BF25, Section 9.2]).
4. The CM case and magnetic modular forms

### v2

@@ Acknowledgements
The author would like to thank many people for useful discussions on this project, including Kilian Bönisch, Steven Charlton, Chi Hong Chow, Henri Cohen, Håvard Damm-Johnsen, Vasily Golyshev, Wen-Ching Winnie Li, Yingkun Li, Ling Long, Alexandre Maksoud, Danylo Radchenko, Berend Ringeling, Fernando Rodriguez Villegas, Hasan Saad, and Jan Vonk. The author is in particular grateful to Don Zagier and Wadim Zudilin for all the useful ideas on proofs and generalizations.
This project originated from discussions during the International Conference on Modular Forms and $q$-Series at University of Cologne, and during the follow-up workshop at the Hausdorff Research Institute for Mathematics to the trimester program “Periods in Number Theory, Algebraic Geometry and Physics” funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. Part of the later discussions also took place during the author’s visit to the Abdus Salam International Centre for Theoretical Physics hosted by Vasily Golyshev and Don Zagier, and during the author’s visit to Louisiana State University hosted by Ling Long and partially supported by the Simons Foundation grant MP-TSM-00002492.
@@ 2. Simple poles
We first recall that Theorem 1.7 implies that for all good primes $p\geq 5$ of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ and all $n\in\mathbb{Z}^{+}$,
$a_{np}(F_{k,C})\;\equiv\;a_{p}(C)^{k-2}a_{n}(F_{k,C})\pmod{p}$.
From a different point of view, this describes the behavior of $F_{k,C}$ under the $U_{p}$ operator. In particular, it says that $F_{k,C}$, when viewed as a mod $p$ modular form, is a $U_{p}$-eigenform with eigenvalue $a_{p}(C)^{k-2}$. In general, one can also describe the behavior of $F_{k,C}$ under the $U_{p}^{l}$ operator if one considers the supersingular and the ordinary cases separately.
@@
While Theorem 1.6 already presents a connection between meromorphic modular forms and hypergeometric functions, the type of congruences that appear in 2.1 and 2.2 may also remind readers of hypergeometric (super)congruences.
In the case of CM elliptic curves, it is possible to unify the supersingular and ordinary dichotomy via a single, but slightly weaker, three-term ASD congruence. There are two viewpoints in the CM setting: one can interpret everything in terms of CM elliptic curves or in terms of negative discriminants. We will elaborate on both viewpoints here.
@@
$a_{np^{l}}(F_{k,C})-a_{p}(\Theta){\hskip 1.0pt}a_{np^{l-1}}(F_{k,C})+\chi_{K}(p)p^{k-2}a_{np^{l-2}}(F_{k,C})\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
@@
By Deuring’s criterion (see [Deu41] or [Sil94, Chapter II, Exercise 2.30]), a good prime $p$ of $C$ is ordinary (resp. supersingular) if and only if $\chi_{K}(p)=1$, i.e., $p$ splits in $K$ (resp. $\chi_{K}(p)=-1$, i.e., $p$ is inert in $K$). The coefficient $a_{p}(\Theta)$ for a good prime $p$ of $C$ can then be written down explicitly as
@@
where $\alpha_{p}$ and $\beta_{p}$ are the two roots of $X^{2}-a_{p}(C)X+p$ when $\chi_{K}(p)=1$.
From this explicit description, one can readily deduce 2.4 from the supersingular (2.1) and ordinary (2.2) dichotomy. In the supersingular case, i.e., $\chi_{K}(p)=-1$, 2.4 reduces to that
@@
which follows from 2.1. In the ordinary case, i.e., $\chi_{K}(p)=1$, 2.4 reduces to that
@@
$\left(a_{np^{l}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\right)$
$\hskip 42.67912pt+\bigg(\frac{p}{u_{p}(C)}\bigg)^{k-2}\left(a_{np^{l-1}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-2}}(F_{k,C})\right)\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
which then follows from 2.2.
Now, we turn to the negative discriminants. Let D<4D<-4 be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $\alpha_{D}\in\mathcal{H}$ be a fixed CM point of discriminant $D$. For $k\in\{4,6,8,10,14\}$, write
@@
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form
@@
One can also write 2.2 in the CM case explicitly in terms of negative discriminants.
@@
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $\big(\frac{D}{p}\big)=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$, and write $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,D})$ $\;\equiv\;\pi^{k-2}a_{np^{l-1}}(F_{k,D})\pmod{\overline{\pi}^{(k-1)l}}$.
@@ 3. Poles of higher order
$\mathrm{MMF}_{k,C}\;:=\;\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}$ (3)
@@
The consideration of this particular space is motivated by several reasons. It is easy to check that for $r\geq k$, $\frac{E_{k}}{(j-j(C))^{r}}$ can be written as a linear combination of $\frac{E_{k}}{j-j(C)},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}$, and $\mathcal{D}^{k-1}f$ for some modular form $f$ of weight 2k2-k, where $\mathcal{D}=q\frac{\mathrm{d}}{\mathrm{d}q}$ is the usual derivative. As the congruence properties of $\mathcal{D}^{k-1}f$ are easy to understand  [^5] From the viewpoint of the magnetic property, $\mathcal{D}^{k-1}f$ is trivially (k1)(k-1)-magnetic, and hence not interesting either., it suffices to consider the denominator only up to the (k1)(k-1)-st power. More importantly, the discussion above yields a non-canonical isomorphism (for $k\in\{4,6,8,10,14\}$ and $j(C)\notin\{0,1728\}$)
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}\;\cong\;\frac{S_{k}^{\mathrm{mero},C}}{\mathcal{D}^{k-1}M_{2-k}^{\mathrm{mero},C}},$ (4)
where $M_{k}^{\mathrm{mero},C}$ denotes the space of meromorphic modular forms of weight $k$ and level $1$ with precisely one pole at (the point on $\mathcal{H}$ corresponding to) $C$, and $S_{k}^{\mathrm{mero},C}$ denotes its subspace consisting of elements with vanishing constant terms in the Fourier expansions. Here a similar quotient where $C$ is replaced by the cusp $\infty$ occurred in Kazalicki–Scholl [KS16], where they interpreted the quotient cohomologically and proved the ASD congruences of weakly holomorphic modular forms. In our case, a cohomological interpretation of the quotient was discussed in Brown–Fonseca [BF25], which was related to $\mathrm{Sym}^{k-2}C$. This suggests that these meromorphic modular forms should be linked to symmetric powers of elliptic curves via the ASD congruences.
The guiding philosophy in our case is that the action of $U_{p}$ on the modular form side  [^6] Indeed, there is no $U_{p}$ action on the space $\mathrm{MMF}_{k,C}$, since the $U_{p}$ operator increases the level and shifts the pole. Still, we will formally consider the $U_{p}$ operator as defined by $\sum a_{n}q^{n}\mapsto\sum a_{pn}q^{n}$. should correspond to the action of the Frobenius at $p$ on the elliptic curve side.
@@
where $P_{p}^{*}$ is the reciprocal polynomial of $P_{p}$, i.e., $P_{p}^{*}(X)=X^{k-1}P_{p}(X^{-1})$.  [^7] In view of Kazalicki–Scholl [KS16], the formulation using $V_{p}$ may be the “correct” one, since their proof of the ASD congruences involves identifying a $V_{p}$-related operator with the Frobenius at $p$.
Remark 3.3.
Note that $(k-1)l-\frac{1}{2}(k-3)k-r$ is never positive when $l=1$ and $r=1$. It is thus not possible to recover Theorem 1.7 from 3.1 by choosing $l=1$ and $r=1$. This is not the only case where the actual congruences appear stronger than the ones predicted by 3.1. For example, let $k=6$ and assume the same conditions as in 3.1. Then, it appears that the following are true: for all $n\in\mathbb{Z}^{+}$ and $1\leq r\leq 5$,
$a_{np^{2}}(F_{6,C}^{(r)})+c_{p,1}^{(6)}{\hskip 1.0pt}a_{np}(F_{6,C}^{(r)})+c_{p,2}^{(6)}{\hskip 1.0pt}a_{n}(F_{6,C}^{(r)})$  $\;\equiv\;0\pmod{p^{4-r}}$
$a_{np^{3}}(F_{6,C}^{(r)})+c_{p,1}^{(6)}{\hskip 1.0pt}a_{np^{2}}(F_{6,C}^{(r)})+c_{p,2}^{(6)}{\hskip 1.0pt}a_{np}(F_{6,C}^{(r)})+c_{p,3}^{(6)}{\hskip 1.0pt}a_{n}(F_{6,C}^{(r)})$  $\;\equiv\;0\pmod{p^{7-r}}$.
We give an example of 3.1 when $k=6$ and $r=2$, which also illustrates the stronger congruences described in Remark 3.3.
Example 3.1.
Let $k=6$, $r=2$, and $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}+y=x^{3}-x^{2}$ with LMFDB label 11.a3 and j(C)=4096/11j(C)=-4096/11. We will look at
$F\;=\;F_{6,C}^{(2)}\;=\;\frac{E_{6}}{(j+4096/11)^{2}}.$
In this case,
$(k-1)l-\tfrac{1}{2}(k-3)k-r\;=\;5l-11.$
Now, let $p=5$, which is a good prime of $C$ and satisfies that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. In this case, we have $a_{5}(C)=1$, and one computes the characteristic polynomial of the Frobenius at $5$ on $\mathrm{Sym}^{4}C$ to be
$P_{5}(X)\;=\;X^{5}-11X^{4}-495X^{3}+12375X^{2}+171875X-5^{10}.$
Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{2}n}(F)-11{\hskip 1.0pt}a_{5n}(F)-495{\hskip 1.0pt}a_{n}(F)$ $\pmod{5^{2}}$
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{3}n}(F)-11{\hskip 1.0pt}a_{5^{2}n}(F)-495{\hskip 1.0pt}a_{5n}(F)+12375{\hskip 1.0pt}a_{n}(F)$ $\pmod{5^{5}}$
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{4}n}(F)-11{\hskip 1.0pt}a_{5^{3}n}(F)-495{\hskip 1.0pt}a_{5^{2}n}(F)+12375{\hskip 1.0pt}a_{5n}(F)+171875{\hskip 1.0pt}a_{n}(F)$ $\pmod{5^{9}}$
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{5}n}(F)-11{\hskip 1.0pt}a_{5^{4}n}(F)-495{\hskip 1.0pt}a_{5^{3}n}(F)+12375{\hskip 1.0pt}a_{5^{2}n}(F)+171875{\hskip 1.0pt}a_{5n}(F)-5^{10}\cdot a_{n}(F)$ $\pmod{5^{15}}$.
Here the congruences when $l=2,3$ are stronger than the ones predicted by 3.1, as discussed in Remark 3.3; the one when $l=4$ is the same as predicted by 3.1; the one when $l=5$ is stronger again, but for a general prime, it does seem that the one when $l=5$ is usually the same as predicted by 3.1.
If one separates the supersingular and ordinary primes as before, then one also observes a generalization of 2.1. However, generalizations of 2.2 seem to only exist in the CM case, which will be discussed in Section 4.
@@
When $p$ is supersingular, $F_{k,C}^{(r)}$ may be viewed as an overconvergent $p$-adic modular form, since the pole lies in the supersingular locus. Indeed, for $l\geq 2$, 3.4 leads to the consideration of $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$. Here one observes that there is no old cusp form by our choice of $k$ and the operator $U_{p}-p^{k-2}$ annihilates all the newforms of weight $k$ and level $p$. Hence, heuristically one should be able to write $(U_{p}-p^{k-2})(F_{k,C}^{(r)})$ as an asymptotic sum of $U_{p}$-eigenforms  [^8] We refer to [Cal, Section 3.5] and [GM95] for discussions on the asymptotic expansions of overconvergent modular forms using $U_{p}$-eigenforms. with slopes $\geq k-1$, which then also supports the conjectural congruence satisfied by $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$.
@@
3.4 implies 3.1 in the supersingular case. We sketch a proof here for $l\geq k-1$. For a supersingular prime $p$, the characteristic polynomial $P_{p}(X)$ is given by
@@
It then follows from 3.4 and direct computations that for all $s\in\mathbb{Z}_{\geq 0}$,
@@
We give an example of 3.4 when $k=4$ and $l=1$, In this case, one expects that
$a_{np}(F_{4,C}^{(1)})\;\equiv\;0\pmod{p^{2}}\quad\text{ and }\quad a_{np}(F_{4,C}^{(2)})\;\equiv\;0\pmod{p}$
for all $n\in\mathbb{Z}^{+}$, or equivalently,
$U_{p}F_{4,C}^{(1)}\;\equiv\;0\pmod{p^{2}}\quad\text{ and }\quad U_{p}F_{4,C}^{(2)}\;\equiv\;0\pmod{p}$.
This example is also related to the heuristic described in Remark 3.5.
Example 3.2.
Let $k=4$. Let $p=13$ and let $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}=x^{3}-x^{2}-4$ with LMFDB label 56.b2. In particular, $C$ is supersingular at $13$ and j(C)=4/7j(C)=-4/7 satisfies that $v_{13}(j(C))=0=v_{13}(j(C)-1728)$. We will now try to write
$\frac{E_{4}}{j+4/7}\pmod{13^{2}}\quad\text{ and }\quad\frac{E_{4}}{(j+4/7)^{2}}\pmod{13}$
in terms of linear combinations of Hecke eigenforms of level $13$ and weight $4$.
Note that there are exactly three Hecke eigenforms in $S_{4}(\Gamma_{0}(13))$: the one $f_{1}$ with rational coefficients and with LMFDB label 13.4.a.a, and two embeddings $f_{2},f_{3}$ of the one with coefficients in $\mathbb{Q}(\sqrt{17})$ and with LMFDB label 13.4.a.b. In particular,
$U_{13}f_{1}\;=\;-13f_{1},\quad U_{13}f_{2}\;=\;13f_{2},\quad U_{13}f_{3}\;=\;13f_{3},$
and
$f_{1}$ $\;\equiv\;q-5q^{2}-7q^{3}+17q^{4}+O(q^{5})$ $\pmod{13^{2}}$
$f_{2}$ $\;\equiv\;q-7q^{2}+25q^{3}-11q^{4}+O(q^{5})$ $\pmod{13^{2}}$
$f_{3}$ $\;\equiv\;q-5q^{2}+19q^{3}-9q^{4}+O(q^{5})$ $\pmod{13^{2}}$.
Moreover, it is easy to check that
$f_{1}\;\equiv\;f_{3}\pmod{13}$.
For $r=1$, we find that
$\frac{E_{4}}{j+4/7}\;\equiv\;q+99q^{2}+32q^{3}+108q^{4}+O(q^{5})\;\equiv\;72f_{1}+117f_{2}+150f_{3}\pmod{13^{2}}$. (5)
Hence,
$U_{13}\bigg(\frac{E_{4}}{j+4/7}\bigg)\;\equiv\;13\cdot(-72f_{1}+117f_{2}+150f_{3})\;\equiv\;13\cdot(78f_{1}+117f_{2})\;\equiv\;0\pmod{13^{2}}$.
For $r=2$, we find that
$\frac{E_{4}}{(j+4/7)^{2}}\;\equiv\;q^{2}+10q^{3}+q^{4}+O(q^{5})\;\equiv\;7f_{1}+6f_{2}\pmod{13}$. (6)
Hence,
$U_{13}\bigg(\frac{E_{4}}{(j+4/7)^{2}}\bigg)\;\equiv\;13\cdot(-7f_{1}+6f_{2})\;\equiv\;0\pmod{13}$.
Here the congruences (5) and (6) are provable by multiplying both sides by $j+4/7$ or $(j+4/7)^{2}$ and comparing the Fourier coefficients up to a suitable number of terms, so we in fact obtain a proof of 3.4 in this special case.
@@
We finish this section with two comments. The first comment is that one should be able to generalize the whole setting to elliptic curves over number fields. In this case, one should then consider congruences modulo powers of some fixed prime ideal $\mathfrak{p}$, and the action of the Frobenius at $\mathfrak{p}$ on the elliptic curve side should correspond to the action of $U_{\mathrm{N}(\mathfrak{p})}$ on the modular form side, where $\mathrm{N}(\mathfrak{p})$ denotes the norm. The second comment is that one should expect a motivic interpretation of the connection between meromorphic modular forms and symmetric powers of elliptic curves  [^9] This is also why we always write $\mathrm{Sym}^{k-2}C$ instead of $\mathrm{Sym}^{k-2}H^{1}(C)$ with some specified cohomology, since we would like to think of $\mathrm{Sym}^{k-2}C$ as being the motive attached to the symmetric power of the elliptic curve., as discussed in Brown–Fonseca [BF25]. In particular, the congruences that we observe here should be viewed as visualizations of the motivic connection at finite places, and it may be reasonable to expect visualizations at infinite places linking special values of $L$-functions of symmetric powers of elliptic curves to “periods” of meromorphic modular forms.
4. The CM case, supercongruences, and magnetic modular forms